Write the equation of the line that passes through point (1,-8) with a slope of 0.
Answer by longjonsilver(2297)

any straight line is of the form y=mx+c.

If gradient is zero then that means m=0. We also know that x=1 when y=-8. So, we get -8=(0)(1)+c
--> c=-8

So, the equation is, starting from y=mx+c:
y=-8

jon